OpenStudy (anonymous):

help with binomial theorem? expand (x+3)^4

OpenStudy (campbell_st):

which part... is causing you problems..?

OpenStudy (anonymous):

the expansion, i missed the lesson at school so i really dont understand

OpenStudy (campbell_st):

ok... so the 4th power... the coefficients come from pascal's triangle  so start with \[1 \times x^{4-0}(3)^{0} + 4 \times(x)^{4 -1}(3)^{1} + 6\times x^{4-2}(3)^{2}+...\] you just need to continue the pattern... into you have x^0 then just evaluate the powers of 3 and multiply by the coefficient... hope this makes sense